OpenStudy (anonymous):

The base of an isosceles triangle is half the length of each of its congruent sides. If the perimeter of the triangle is 120 inches, what is the length of the base? A. 12 inches B. 24 inches C. 16 inches D. 6 inches

OpenStudy (anonymous):

please help i am really stuck on this question

OpenStudy (jchick):

Isosceles triangles in particular, the base angles have the same measure Or A = C

OpenStudy (jchick):

A = \frac{ bh _{b} }{ ? }}\]

OpenStudy (jchick):

\[\frac{ bh _{b}}{ 2 }\]

OpenStudy (jchick):

Here's your triangle. x = base 2x = each of the congruent sides. Add all the sides up to equal 120. x + 2x + 2x = 120 Simplify. 5x = 120 x = 24 Base would = 24 inches.

OpenStudy (anonymous):

thanks :D

OpenStudy (jchick):

No problem!